### v1

# Engineering dipole-dipole couplings for enhanced cooperative light-matter interactions

Abstract
Cooperative optical effects are enabled and controlled by interactions between molecular dipoles, meaning that their mutual orientation is of paramount importance to, for example, superabsorbing light-harvesting antennas. Here we show how to move beyond the possibilities of simple geometric tailoring, demonstrating how a metallic sphere placed within a ring of parallel dipoles engineers an effective Hamiltonian that generates “guide-sliding” states within the ring system. This allows steady-state superabsorption in noisy room temperature environments, outperforming previous designs while being significantly simpler to implement. As exemplified by this showcase, our approach represents a powerful design paradigm for tailoring cooperative light-matter effects in molecular structures that extends beyond superabsorbing systems, to a huge array of quantum energy transport systems.
Dipolar interactions underpin countless important physical [1, 2, 3], chemical [4, 5] and biological phenomena [6, 7], with their influence extending to a wide variety of photophysics [8, 9]. Acting between molecules possessing permanent or induced dipole moments, in their simplest form, they originate from Coulomb interactions of the charged constituents of each (overall neutral) molecule, exhibiting rich orientation-dependence. This leads to notably different couplings in structures whose molecules are aligned ‘face-to-face’ (known as H-aggregates), distinct from those in the ‘head-to-tail’ configuration (J-aggregates) [8].
One prominent contemporary use of the dipole-dipole interaction is the enhancement of effective optical absorption rates, as motivated by Dicke’s prediction of superradiant atomic dipoles [10, 11, 12, 13, 14, 15]. The concept of the time reversed process — superabsorption — achievable by manipulating symmetric dipolar interactions to sustain a superabsorbing state has been proposed [16, 9], and experimentally observed in scenarios circumventing the need for dipolar interactions [18, 19]. The concept draws inspiration from photosynthetic systems where J-aggregate-like ring arrangement reduces light-matter coupling due to destructive interference of the transition dipoles in the plane of the ring and can prevent damage by limiting solar absorption [20] and efficient long-range exciton transport through dark states [21]. Conversely, artificial systems arranged as H-aggregates promise enhanced light-harvesting performance for improved solar energy conversion [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32].
A question then naturally presents itself — is it possible to have the ‘best of both worlds’, where, for example, an H aggregate with desirable far-field properties is imbued with near-field couplings associated with a J aggregate? Here we answer this question positively by exploiting the fact that the textbook dipole-dipole interaction formula is not an immutable law of nature but is instead a particular case of a broader class dipole-dipole interactions in the presence of dielectric or metallic environments [33, 34]. This is the realm of macroscopic QED [35, 36], where techniques originating in the study of the Casimir [37] and Casimir-Polder [38] effects (examples of dispersion forces [1, 40]) can be used to engineer interatomic interaction processes almost at will, most recently in the rapidly-expanding field of inverse design [41].
[…]
Figure 1: (a) A rendering of the parallel guide-slide super absorber, placing a dielectric sphere in the centre of the ring engineers the dipole-dipole couplings. (b) In turn these lead to a level structure that ‘guides’ excitations onto optically brighter (approximately Dicke-ladder) states via phonon relaxation. (c) Parallel dipoles exhibit $N^{2}$ enhancement of the transition dipole strength near Dicke-ladder half-inversion, allowing for superabsorption.
The free space coupling $J_{ij}^{(0)}$ between two dipoles $\mathbf{d}_{i}$ and $\mathbf{d}_{j}$ at positions $\mathbf{r}_{i}$ and $\mathbf{r}_{j}$ separated by $\bm{R}=\mathbf{r}_{i}-\mathbf{r}_{j}$ is
$J_{ij}^{(0)}=\frac{1}{4\pi|\bm{R}|^{3}}\left[\mathbf{d}_{i}\cdot\mathbf{d}_{j}-3(\mathbf{d}_{i}\cdot\mathbf{e}_{R})(\mathbf{d}_{j}\cdot\mathbf{e}_{R}\right)],$ (1)
where $\mathbf{e}_{R}=\bm{R}/|\bm{R}|$ and we have used a system of natural units with $c=\hbar=\varepsilon_{0}=1$. The QED generalisation of Eq. (1) to an arbitrary environment is [42, 43]:
$J_{ij}=-\omega^{2}\mathbf{d}_{i}\cdot\mathrm{Re}[\mathbb{G}(\mathbf{r}_{i},\mathbf{r}_{j},\omega)]\cdot\mathbf{d}_{j},$ (2)
where $\omega$ is the average of the transition frequencies of the two dipoles (here assumed identical), and $\mathbb{G}$ is the dyadic Green’s tensor that solves the Helmholtz equation with boundary conditions determined by the geometry at hand. The choice of a particular geometry and material defines $\mathbb{G}$. For instance, choosing vacuum recovers Eq. (1), as demonstrated in the Supplementary Material (SM), offering flexible design of dipole-dipole interactions through Eq. (2), even for dipoles with fixed relative orientation.
In this article, we showcase the power of this approach through a particular example: employing engineered dipole-dipole interactions enables a so-called ‘guide-slide’ superabsorber that achieves enhanced absorption courtesy of a carefully calibrated interplay between its vibrational and optical environments [9]. This crucially relies on an effective structuring of energy levels due to Lamb-shifts induced by a dielectric metal sphere placed in the centre of a molecular dipole ring (see Fig. 1) — a simple type of inverse design. Achieving this without such engineered interactions necessitates sub-optimal alignment of dipoles (with reduced collective dipole moment—counterproductive to superabsorption) in addition to ad-hoc re-initialisation, and placement of the structure in a photonic band gap [9]. As we presently argue, engineering of the dipole-dipole interaction reduces the importance of such external control, enabling the design of simpler yet more powerful photonic devices.
We begin by illustrating the engineering of dipolar interactions with the minimal system for analytically demonstrating the tunable conversion of a physical H aggregate into an effective J aggregate. We consider two dipoles in vacuum in the region $z>0$, with the region $z<0$ filled with a material of relative permittivity $\varepsilon$ (and unit relative permeability), in the electrostatic limit where any separations are assumed much shorter than the (reduced) transition wavelength $\lambda/2\pi$. The Green’s tensor for this situation is particularly simple (see [2, 3]) and is shown in the SM. Substituting that expression into Eq. (2) produces an analytic result for the ratio of the interaction matrix element near a planar surface to its free space value ${J^{(0)}_{ij}}$. Specialising to a configuration of interest, namely dipoles at the same distance $z$ from the half-space, both with dipole moments oriented in the $y$ direction and separated in the $x$ plane by a distance $\Delta x$ (as illustrated in Fig. 2), one finds ${J_{ij}}/{J^{(0)}_{ij}}=1-\frac{\varepsilon-1}{\varepsilon+1}[(2z/\Delta x)^{2}+1]^{-3/2}$ (see SM). Since we seek cases where the interaction matrix element changes sign relative to free space, we require combinations of $z/\Delta x$ and $\varepsilon$ for which ${J_{ij}}/{J^{(0)}_{ij}}<0$, achievable for $\varepsilon<-1$ (implying that a surface plasmon is the crucial ingredient) and distance to wall $z$ sufficiently shorter than pair separation $\Delta x$, as shown in Fig. 2. Under these conditions, an H-aggregate close to a surface supporting a surface plasmon assumes the interaction matrix element of a J-aggregate, as if its dipoles had been rotated.
Figure 2: (a) Sign of the interaction matrix element $J_{ij}$ as the permittivity and dipole positions are varied near a dielectric surface as indicated in (b). The shaded region indicates where the sign of the interaction matrix element is flipped relative to the same aggregate in free space (matrix element $J_{ij}^{(0)}$). The blue dashed line shows where the interaction matrix element equals that of a J-type aggregate (dipoles placed end-to-end) in free space with the same separation $\Delta x$.
Building on the dimer, we proceed to a ring of $N\geq 3$ dipoles, where a changed coupling sign is key to supporting superabsorbing behaviour [9]. Previously, this was accomplished in a ring arrangement with dipole moments tilted appropriately with respect to each other. By contrast, engineering the dipolar interactions will allow us to maximise the collective far-field absorption strength, while achieving the requisite energetic ordering through negative couplings with the help of a metallic sphere inside the ring as illustrated in Fig. 1. These properties lead to the generation of a guide-slide superabsorber.
We define a guide-slide superabsorber to be a collection of optical dipoles with properties: (1) a ladder of excitation manifolds of a fixed number of dipole excitations, each with rapid relaxation to a well-defined lowest-energy state within the manifold, (2) enhanced optical rates due to collective effects between dipoles coupling the lowest-energy states of adjacent manifolds [9]. To show the efficacy of our surface-plasmon-enabled guide-slide system we will utilise an open quantum systems heat engine framework and compare our new design against the previously proposed tilted guide-slide antenna [9].
We consider a ring assembly of interacting chromophores, each approximated as a two-level optical dipole. All dipoles couple to the common electromagnetic field and to individual vibrational baths, yielding the total Hamiltonian
$H=H_{S}+H_{\text{opt}}+H_{I,\text{opt}}+H_{\text{vib}}+H_{I,\text{vib}},$ (3)
where the system Hamiltonian describes the energies associated with just the dipoles
$H_{S}=\omega_{0}\sum_{i=1}^{N}\sigma_{+}^{(i)}\sigma_{-}^{(i)}+\sum_{i\neq j=1,}^{N}J_{ij}\sigma_{+}^{(i)}\sigma_{-}^{(j)}$ (4)
with bare transition energies $\omega_{0}$ and $J_{ij}$’s according to Eq. (2); the $\sigma_{\pm}^{(n)}$ operators are the raising and lowering operators for the $n$th dipole. The free optical field obeys $H_{\text{opt}}=\sum_{p}\omega_{p}a_{p}^{\dagger}a_{p},$ with $\omega_{p}$ and $a^{({\dagger})}_{p}$, representing the energy and annihilation (creation) operator, respectively, associated with the $p$th electromagnetic field mode. Dipole-electromagnetic interactions are captured by $H_{I,\text{opt}}=\sum_{i=1}^{N}\sigma_{x}^{(i)}\sum_{p}\mathbf{d}_{i}\cdot\mathbf{e}_{p}f_{p}(a_{p}+a_{p}^{\dagger}),$ where $f_{p}$ determines the coupling strength of the $p$th photon mode (polarisation vector $\mathbf{e}_{p}$) to the dipole with vector dipole $\mathbf{d}_{i}$. Finally, each dipole couples to its own vibrational environment, with interaction Hamiltonian $H_{I,\text{vib}}=\sum_{i=1}^{N}\sigma_{z}^{(i)}\sum_{v}g_{i,v}(b_{i,v}+b^{\dagger}_{i,v})$ and $H_{\text{vib }}=\sum_{i=1}^{N}\sum_{v}\omega_{i,v}b_{i,v}^{\dagger}b_{i,v},$ where $\omega_{i,v}$, $g_{i,v}$, and $b_{i,v}^{({\dagger})}$ represent, respectively, the energy, coupling strength, and annihilation (creation) operator of mode $v$ at the $i$th dipole. Geometrically, we fix a nearest neighbour dipole separation of 2.5 nm as a natural length scale for molecular chromophore separations [46, 47], and sufficient for placing a metallic nanoparticle within the ring. The dipoles are approximately spatially indistinguishable by the electromagnetic field for relevant optical frequencies ($\approx 700/2\pi\,$nm2.5much-greater-thanabsent2.5$\gg$ 2.5≫ 2.5 nm) due to slowly varying phases across the ring. Furthermore, we let the radius of the inserted sphere be 1 nm shorter than the ring radius.
The specific eigenstructure of our system Hamiltonian $H_{S}$ (4) will determine how optical and vibrational baths give rise to dissipative energy flow and conversion within an open quantum systems framework [22, 48, 25, 28]. We expect maximal interaction with the optical field for dipoles aligned perpendicular to the ring plane, as individual dipoles become indistinguishable and form an effective super-dipole. Without the sphere, such a configuration in free space would result in an H-aggregate system with positive-valued $J_{ij}$ coupling elements. However, H-aggregates feature optically bright states at higher energy than optically dark states. Vibrational relaxation would result in rapid population transfer from the bright to the dark states, effectively trapping excitations in optically inactive states (see Fig. 3a), preventing superradiance and superabsorption.
To mitigate against such phonon-induced breakdown of superabsorption, a tilted guide-slide configuration was proposed [9], wherein each of the dipoles is tilted sufficiently from the azimuth to achieve a negative nearest neighbour interaction (in turn positioning dark states energetically above bright states). Then vibrational relaxation funnels excitations onto the Dicke ladder. However, the imperfect alignment of the dipoles also reduces the collectively enhanced dipole perpendicular to the ring plane, shifting some transition oscillator strength into destructive interference in the plane.
Strategic engineering of our nearest neighbour coupling, by the metallic sphere placed within the dipole ring, generates both negative nearest neighbour couplings and perfectly aligned dipoles in concert. This enables a guide-sliding effect due to vibrational relaxation alongside optimally enhanced optical rates courtesy of fully aligned dipoles. The level scheme depiction in Fig. 3b shows that the most optically active states lie at the bottom of each manifold. With vibrational relaxation occurring much faster than optical decay, this setup fulfils our guide-slide conditions outlined above.
[…]
Figure 3: Level transition schematic of a four-dipole ring (a) in free space (b) as spherical guide-slide superabsorber. Blue dotted arrows represent fast phonon relaxation inside manifolds of fixed excitation number, red arrows represent optical transitions between manifolds. Arrow saturation is scaled by the relative strength of the transition. Intra-manifold energy splitting is accentuated relative to manifold spacings.
We model the dynamics of our ring antenna within the Born-Markov approximation leading to a Bloch-Redfield master equation [7]. This generates non-unitary dynamics for the reduced density matrix of the dipole ring due to dissipative influences from the vibrational and photon environments. The approximations implicit within this model are that we are utilising second-order perturbation theory for the system-bath coupling and that the bath relaxation timescale is much faster than the dynamics associated with the system, such that the bath correlations can be considered as delta functions, see SM for more details. This results in a master equation of the form $\dot{\rho}_{S}(t)=-i[H_{S},\rho_{S}]+(\mathcal{D}_{\text{opt}}+\mathcal{D}_{\text{vib}})\rho_{S}$.  The optical dissipator $\mathcal{D}_{\text{opt}}$ is responsible for transitions between excitation number manifolds, through which sunlight photons pump our superabsorber. Conversely, the vibrational dissipator, $\mathcal{D}_{\text{vib}}$, leads to relaxation within excitation manifolds, and, dependent on eigenstructuring, leads to either guide-sliding or dark-state trapping.
To quantify the efficiency of the guide-slide superabsorber, we employ a trap model for extracting and converting ring excitons to useful energy [50, 51, 52]. This is modelled by incoherently coupling an additional two-level system to the ring as an abstract load [50, 48], implemented through an additional dissipator of Lindblad form in the master equation $\mathcal{D}_{X}\rho_{S}=\gamma_{X}(X\rho_{S}X^{\dagger}-\frac{1}{2}\{X^{\dagger}X,\rho_{S}\}),$ where $\gamma_{X}$ is the rate of extraction from the ring to the trap and the Lindblad operator $X$ defining this incoherent transport is $X=\sum_{i=1}^{N}\sigma_{-}^{(i)}\otimes\sigma_{+}^{(t)}.$ Such an extraction process can be interpreted by treating the trap as another dipole displaced below the centre of the ring and considering the second-order effects from its coupling through the photon field, e.g. as a FRET process.
Additionally, we introduce a trap decay dissipator, causing incoherent transitions from the excited state of the trap to its ground state of the form $\mathcal{D}_{t}\rho_{S}=\Gamma_{t}(\sigma_{-}^{(t)}\rho_{S}\sigma_{+}^{(t)}-\frac{1}{2}\{\sigma_{+}^{(t)}\sigma_{-}^{(t)},\rho_{S}\}).$ We then construct the total master equation for the reduced density matrix by summation of these dissipators
$\frac{d}{dt}\rho_{S}=-i[H_{S},\rho_{S}]+(\mathcal{D}_{\text{opt}}+\mathcal{D}_{\text{vib}}+\mathcal{D}_{X}+\mathcal{D}_{t})\rho_{S}.$ (5)
Combining high-temperature solar photons with room-temperature phonons and the trap, we have constructed an effective quantum heat engine (QHE), allowing us to determine the power output of our ring antennas, and uncover how this power scales with the number of dipoles. To this end, we calculate the current of the QHE given by $I=e\Gamma_{t}\langle\rho_{t,e}\rangle,$ where $\langle\rho_{t,e}\rangle$ is the expectation value of the steady state population for the excited state of the trap. Using this formalism we also calculate the potential difference generated as given by $eV=\hbar\omega_{t}+k_{B}T_{\text{vib}}\ln{\left[\frac{\langle\rho_{t,e}\rangle}{\langle\rho_{t,g}\rangle}\right]},$ where $T_{\text{vib}}$ is the temperature of the vibrational baths and $\langle\rho_{t,g}\rangle$ is the expectation of the steady state population for the trap’s ground state [50]. From this, we compute the power output of the solar absorber at the steady state by $P=IV$. Steady states are calculated by finding the null-space of Eq. (5). By allowing for $\Gamma_{t}$ to vary, we may find the optimal power output for each configuration of the ring system, such that $P_{\text{max}}=\text{max}_{\Gamma_{t}}\{I(\Gamma_{t})V(\Gamma_{t})\}$.
Utilising this framework, we directly compare the power outputs of the solar absorbers using the tilted guide-slide developed previously and the novel surface plasmon-enabled (parallel) guide-slide developed here. Fig. 4 shows the results of numerical simulations for varying numbers of dipoles in the rings, for a permittivity of $\varepsilon=-2.37$, similar to that of chromium at $1.8$eV. Fig. 4(c) depicts the power output per dipole, and we note an increase with $N$ for the spherical guide-slide setup, showing super-extensive scaling. Additionally, we see for small values of $N$ the power output achieved by the tilted guide-slide is greater than that of the parallel guide-slide. This is due to the relative size of the sphere to the ring radius being small in this limit, restricting the ability to engineer dipolar interactions. By contrast, increasing ring radius improves the approximation of dipoles experiencing a planar dielectric interface, meaning the design is scalable to large $N$. In Fig. 4(b), a log-log plot shows the exponent for each configuration’s power output of the form $P=\alpha N^{m}$. Our surface plasmon-enabled guide-slide has greater super-extensive power scaling than the tilted guide-slide, with a growth exponent $m=1.55$ vs. $m=1.08$. This increase in super-linear scaling is due to the parallel dipoles in the surface plasmon-enabled guide-slide offering maximal collective enhancement.
Our results exemplify the general utility of selectively designing the local electromagnetic field through metals or dielectrics to achieve bespoke couplings different from far-field behaviour. While incorporating a spherical metallic nano-particle into the centre of a nanoscopic molecular ring of dipoles presents an implementation challenge, recent experimental studies have validated the approach of utilising DNA origami to interface molecular rings with nano-particles  [53, 54, 55], supporting the in-principle feasibility of our outlined design. Its simplicity, enhanced super-linear scaling, and operability without a photonically engineered environment and active re-initialisation establish our present design as a leading candidate for a quantum-enhanced molecular antenna. As we elaborate in the SM, structuring the local electromagnetic field modes to induce the engineered dipole-dipole interactions also impacts the optical decay rates and yields a secondary advantage. Specifically, we observe an additional enhancement of the super-linear behaviour from a $N$-dependent Purcell effect. However, as this effect is not directly related to realising the effect of superabsorption as the counterpart to Dicke superradiance, we have opted not to include this effect in the discussion here to avoid ambiguity. Further, while our design—unlike previous superabsorbing ring proposals—performs well without a photonic band-gap and active reinitisalition, it benefits and maintains its advantage when available (see SM).
[…]
Figure 4: Scaling of the solar cell power output for the configuration of the parallel spherical ring guide-slide, and the previously developed tilted guide-slide. The (a) power output (b) log-log plot of the power output (c) power output per dipole of the ring for increasing number of dipoles in the ring. The super-linear scaling for each configuration $P=\alpha N^{m}$ was calculated. The parallel sphere guide-slide has $m=1.55$ and the tilted guide-slide $m=1.08$. The theoretical maximum Dicke scaling of $N^{2}$ (green dashed) and linear scaling (black dashed) are also plotted for comparison in the log-log graph. Full simulation parameters are given in the SM.
In this article, we have demonstrated that surface plasmons can be exploited to engineer molecular dipole-dipole interactions, thereby enabling advantageous couplings for light-harvesting systems. Building upon this, we have developed a novel superabsorber by precisely engineering the Hamiltonian of a molecular dipole ring through the strategic placement of a metallic nanoparticle at its center. Within the framework of a quantum heat engine, we have shown that our design surpasses previous proposals in performance, and, due to its greater simplicity, represents an ideal candidate for the development of light-harvesting superabsorbers. Importantly, we have illustrated that the innovative design of the local dielectric or metallic environment allows for the engineering of system Hamiltonians with properties conducive to the design and implementation of quantum-enhanced devices, representing a critical advancement towards realising efficient energy transport systems.
Acknowledgments
The authors thank Will Brown for helpful discussions. A. Burgess and E.M. Gauger thank the Leverhulme Trust for support through grant number RPG-2022-335. M.C. Waller thanks the EPSRC Doctoral Training Programme EPSRC/DTP 2020/21/EP/T517896/1, and R. Bennett acknowledges financial support from EPSRC grant EP/W016486/1.