### v2

# Engineering dipole-dipole couplings for enhanced cooperative light-matter interactions

Abstract
Cooperative optical effects are enabled and controlled by interactions between molecular dipoles, meaning that their mutual orientation is of paramount importance to, for example, superabsorbing light-harvesting antennas. Here we show how to move beyond the possibilities of simple geometric tailoring, demonstrating how a metallic sphere placed within a ring of parallel dipoles engineers an effective Hamiltonian that generates “guide-sliding” states within the ring system. This allows steady-state superabsorption in noisy room temperature environments, outperforming previous designs while being significantly simpler to implement. As exemplified by this showcase, our approach represents a powerful design paradigm for tailoring cooperative light-matter effects in molecular structures that extends beyond superabsorbing systems, to a huge array of quantum energy transport systems.
Dipolar interactions underpin countless important physical [1, 2, 3], chemical [4, 5] and biological phenomena [6, 7, 8, 9], with their influence extending to a wide variety of photophysics [10, 11]. Acting between molecules possessing dipole moments they originate from Coulomb interactions of the charged constituents of each (overall neutral) molecule, exhibiting rich orientation-dependence. This leads to notably different couplings in structures whose molecules are aligned ‘face-to-face’ (known as H-aggregates), distinct from those in the ‘head-to-tail’ configuration (J-aggregates) [10].
One prominent contemporary use of the dipole-dipole interaction is the enhancement of effective optical absorption rates, as motivated by Dicke’s prediction of superradiant atomic dipoles [12, 13, 14, 15, 16, 17]. The concept of the time-reversed process — superabsorption — achievable by manipulating symmetric dipolar interactions to sustain a superabsorbing state has recently gained considerable interest in the field theoretically  [18, 9] and circumventing the need for dipolar interactions also experimentally [20, 21]. The concept draws inspiration from photosynthetic systems where J-aggregate-like ring arrangement reduces light-matter coupling due to destructive interference of the transition dipoles in the plane of the ring and can prevent damage by limiting solar absorption [22] and facilitate efficient long-range exciton transport through dark states [23]. Conversely, artificial systems arranged as H-aggregates promise enhanced light-harvesting performance for improved solar energy conversion [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34].
A question then naturally presents itself — is it possible to have the ‘best of both worlds’, where, for example, an H aggregate with desirable far-field properties is imbued with near-field couplings associated with a J aggregate? Here we answer this question positively, exploiting that textbook dipolar interactions are not immutable laws of nature but are instead a particular case of a broader class of interactions in dielectric or metallic environments [35, 36]. This is the realm of macroscopic QED [37, 38], where techniques originating in the study of dispersion forces [39, 40, 1, 42] can be used to engineer interatomic interaction processes almost at will, most recently in the rapidly-expanding field of inverse design [43]. We then exploit these engineered dipole-dipole couplings and show that they enable the design of a high-performance light-harvesting superabsorber, providing a unique synthesis of macroscopic QED and quantum heat-engine formalisms.
[…]
Figure 1: (a) A rendering of the parallel guide-slide super absorber, the placement of a dielectric sphere in the centre of the ring engineers the dipole-dipole couplings. (b) In turn these lead to a level structure that ‘guides’ excitations onto optically brighter (approximately Dicke-ladder) states via phonon relaxation. (c) Parallel dipoles exhibit $N^{2}$ enhancement of the transition dipole strength near Dicke-ladder half-inversion $N_{m}=\lfloor\frac{N}{2}\rfloor$, allowing for superabsorption.
The electrostatic limit of the free space coupling $J_{ij}^{(0)}$ between two dipoles $\mathbf{d}_{i}$ and $\mathbf{d}_{j}$ at positions $\mathbf{r}_{i}$ and $\mathbf{r}_{j}$ separated by $\bm{R}=\mathbf{r}_{i}-\mathbf{r}_{j}$ is
$J_{ij}^{(0)}=\frac{1}{4\pi|\bm{R}|^{3}}\left[\mathbf{d}_{i}\cdot\mathbf{d}_{j}-3(\mathbf{d}_{i}\cdot\mathbf{e}_{R})(\mathbf{d}_{j}\cdot\mathbf{e}_{R}\right)],$ (1)
where $\mathbf{e}_{R}=\bm{R}/|\bm{R}|$ and we have used a system of natural units with $c=\hbar=\varepsilon_{0}=1$. The generalisation of Eq. (1) to an arbitrary environment and separations is [44, 45]:
$J_{ij}=-\omega^{2}\mathbf{d}_{i}\cdot\mathrm{Re}[\mathbb{G}(\mathbf{r}_{i},\mathbf{r}_{j},\omega)]\cdot\mathbf{d}_{j},$ (2)
where $\omega$ is the average of the transition frequencies of the two dipoles (here assumed identical), and $\mathbb{G}$ is the dyadic Green’s tensor that solves the Helmholtz equation with boundary conditions determined by the geometry at hand. The choice of a particular geometry and material defines $\mathbb{G}$. For instance, choosing vacuum recovers Eq. (1), as demonstrated in the Supplementary Material (SM), offering flexible design of dipole-dipole interactions through Eq. (2), even for dipoles with fixed relative orientation.
Here we showcase the power of this approach through a particular example: engineered dipole-dipole interactions enable a ‘guide-slide’ superabsorber that achieves enhanced absorption courtesy of a carefully calibrated interplay between its vibrational and optical environments [9]. This crucially relies on structuring the energy levels due to Lamb-shifts induced by a dielectric metal sphere placed in the centre of a molecular dipole ring (see Fig. 1) Achieving this without engineered interactions necessitates sub-optimal alignment of dipoles (with reduced collective dipole moment—counterproductive to superabsorption) in addition to ad-hoc re-initialisation, and embedding in a photonic band gap [9]. As we presently argue, engineering of the dipole-dipole interaction eliminates the need for such external control, enabling the design of simpler yet more powerful photonic devices, exemplifying the importance of the approach developed.
We begin by illustrating the engineering of dipolar interactions with the minimal system for analytically demonstrating the tunable conversion of a physical H aggregate into an effective J aggregate. This opens up new avenues of research in biophysics and photophysics that depend on dipole-dipole interactions — in our case enabling the design of a new type of light-harvesting superabsorber. We consider two dipoles in vacuum in the region $z>0$, with the region $z<0$ filled with a material of relative permittivity $\varepsilon$ (and unit relative permeability), in the electrostatic limit where any separations are assumed much shorter than the (reduced) transition wavelength $\lambda/2\pi$. [2, 3]) . Substituting the relevant $\mathbb{G}$ ([2, 3]) into Eq. (2) produces an analytic result for the ratio of the interaction matrix element near a planar surface to its free space value ${J^{(0)}_{ij}}$ (see SM). Specialising to dipoles at the same distance $z$ from the half-space, both with dipole moments oriented in the $y$ direction and separated in the $x$ plane by a distance $\Delta x$ ( see Fig. 2), one finds ${J_{ij}}/{J^{(0)}_{ij}}=1-\frac{\varepsilon-1}{\varepsilon+1}[(2z/\Delta x)^{2}+1]^{-3/2}$ (see SM). Since we seek cases where the interaction matrix element changes sign relative to free space, we require combinations of $z/\Delta x$ and $\varepsilon$ for which ${J_{ij}}/{J^{(0)}_{ij}}<0$, achievable for $\varepsilon<-1$ (implying that a surface plasmon is the crucial ingredient) and distance to wall $z$ sufficiently shorter than pair separation $\Delta x$, as shown in Fig. 2. Under these conditions, an H-aggregate close to a surface supporting a surface plasmon assumes the interaction matrix element of a J-aggregate, as if its dipoles had been rotated (see SM for an interpretation via image dipoles). This is a remarkable result that shows for the first time the power of engineering dipole-dipole interactions for improving light harvesting efficiency. This has a wide-reaching impact on the study and development of quantum energy transport and storage systems that rely heavily on such dipolar couplings.
Figure 2: (a) Sign of the interaction matrix element $J_{ij}$ as the permittivity and dipole positions are varied near a dielectric surface as indicated in (b). The shaded region indicates where the sign of the interaction matrix element is flipped relative to the same aggregate in free space (matrix element $J_{ij}^{(0)}$). The blue dashed line shows where the interaction matrix element equals that of a J-type aggregate (dipoles placed end-to-end) in free space with the same separation $\Delta x$.
Building on this simple yet remarkable observation for the dimer, we proceed to a ring of $N\geq 3$ dipoles, where a changed coupling sign is key to supporting superabsorbing behaviour [9]. Previously, this was accomplished in a ring arrangement with dipole moments tilted appropriately with respect to each other. By contrast, engineering the dipolar interactions allows us to maximise the collective far-field absorption strength, while achieving the requisite energetic ordering through negative couplings with the help of a metallic sphere inside the ring as illustrated in Fig. 1.
We define a guide-slide superabsorber to be a collection of optical dipoles with properties: (1) a ladder of excitation manifolds of a fixed number of dipole excitations, each with rapid relaxation to a well-defined lowest-energy state within the manifold, (2) enhanced optical rates due to collective effects between dipoles coupling the lowest-energy states of adjacent manifolds [9]. To show the efficacy of our surface-plasmon-enabled guide-slide system we will utilise an open quantum systems heat engine framework and compare our new design against the previously proposed tilted guide-slide antenna [9].
We consider a ring assembly of interacting chromophores, each approximated as a two-level optical dipole. All dipoles couple to the common electromagnetic field and to individual vibrational baths, yielding the total Hamiltonian
$H=H_{S}+H_{\text{opt}}+H_{I,\text{opt}}+H_{\text{vib}}+H_{I,\text{vib}},$ (3)
where the Coulomb gauge system Hamiltonian describes the energies associated with just the dipoles [33, 48]
$H_{S}=\omega_{0}\sum_{i=1}^{N}\sigma_{+}^{(i)}\sigma_{-}^{(i)}+\sum_{i\neq j=1,}^{N}J_{ij}\sigma_{+}^{(i)}\sigma_{-}^{(j)}$ (4)
with bare transition energies $\omega_{0}$ and $J_{ij}$’s according to Eq. (2); the $\sigma_{\pm}^{(n)}$ operators are the raising and lowering operators for the $n$th dipole. The free optical field obeys $H_{\text{opt}}=\sum_{p}\omega_{p}a_{p}^{\dagger}a_{p},$ with $\omega_{p}$ and $a^{({\dagger})}_{p}$, representing the energy and annihilation (creation) operator, respectively, associated with the $p$th electromagnetic field mode. Dipole-electromagnetic interactions are captured by $H_{I,\text{opt}}=\sum_{i=1}^{N}\sigma_{x}^{(i)}\sum_{p}\mathbf{d}_{i}\cdot\mathbf{e}_{p}f_{p}(a_{p}+a_{p}^{\dagger}),$ where $f_{p}$ determines the coupling strength of the $p$th photon mode (polarisation vector $\mathbf{e}_{p}$) to the dipole with vector dipole $\mathbf{d}_{i}$. Finally, each dipole couples to its own vibrational environment, with interaction Hamiltonian $H_{I,\text{vib}}=\sum_{i=1}^{N}\sigma_{z}^{(i)}\sum_{v}g_{i,v}(b_{i,v}+b^{\dagger}_{i,v})$ and $H_{\text{vib }}=\sum_{i=1}^{N}\sum_{v}\omega_{i,v}b_{i,v}^{\dagger}b_{i,v},$ where $\omega_{i,v}$, $g_{i,v}$, and $b_{i,v}^{({\dagger})}$ represent, respectively, the energy, coupling strength, and annihilation (creation) operator of mode $v$ at the $i$th dipole. Geometrically, we fix a nearest neighbour dipole separation of 2.5 nm inspired by natural molecular chromophore separations [49, 50] and leaving space for a central metallic nanoparticle . Slowly varying phases across the ring render the dipoles approximately spatially indistinguishable by the electromagnetic field for relevant optical frequencies ($\approx 700/2\pi\,$nm2.5much-greater-thanabsent2.5$\gg$ 2.5≫ 2.5 nm). We let the radius of the inserted sphere be 1 nm shorter than the ring radius.
The eigenstructure of our system Hamiltonian $H_{S}$ (4) determines how optical and vibrational baths give rise to dissipative energy flow and conversion within an open quantum systems framework [24, 51, 27, 30]. We expect maximal interaction with the optical field for dipoles aligned perpendicular to the ring plane, as individual dipoles become indistinguishable and form an effective super-dipole. Without the sphere, such a configuration in free space would result in an H-aggregate system with positive-valued $J_{ij}$ coupling elements. H-aggregates feature optically bright states at higher energy than optically dark states. Vibrational relaxation would result in rapid population transfer from the bright to the dark states, effectively trapping excitations in optically inactive states (see Fig. 3a), preventing superradiance and superabsorption.
To mitigate against such phonon-induced breakdown of superabsorption, a tilted guide-slide configuration was proposed [9], wherein each of the dipoles is tilted sufficiently from the azimuth to achieve a negative nearest neighbour interaction (in turn positioning dark states energetically above bright states). Then vibrational relaxation funnels excitations onto the Dicke ladder. However, the imperfect alignment of the dipoles also reduces the collectively enhanced dipole perpendicular to the ring plane, shifting some transition oscillator strength into destructive interference in the plane.
Strategic engineering of nearest neighbour coupling, by the metallic sphere placed within the dipole ring, generates both negative nearest neighbour couplings and perfectly aligned dipoles in concert. This enables a guide-sliding effect due to vibrational relaxation alongside optimally enhanced optical rates courtesy of fully aligned dipoles. Fig. 3b shows that the most optically active states lie at the bottom of each manifold. With vibrational relaxation occurring much faster than optical decay, this setup fulfils our guide-slide conditions outlined above.
[…]
Figure 3: Level transition schematic of a four-dipole ring (a) in free space (b) as spherical guide-slide superabsorber. Blue dotted arrows represent fast phonon relaxation inside manifolds of fixed excitation number, red arrows represent optical transitions between manifolds with manifold number $N_{m}$. Arrow saturation is scaled by the relative strength of the transition. Intra-manifold energy splitting is accentuated relative to manifold spacings. These are calculated by diagonalising $H_{S}$ and calculating transition rates from Fermi’s golden rule.
We model the dynamics of our system with the Bloch-Redfield master equation [7] generating non-unitary dynamics for the reduced density matrix of the dipole ring due to dissipative influences from vibrational and photon environments. The approximations implicit within this model are weak coupling for the system-bath coupling and that the bath relaxation timescale is much faster than the dynamics associated with the system, such that bath correlations can be considered as delta functions, see SM for more details. This results in a master equation of the form $\dot{\rho}_{S}(t)=-i[H_{S},\rho_{S}]+(\mathcal{D}_{\text{opt}}+\mathcal{D}_{\text{vib}})\rho_{S}$.  The optical dissipator $\mathcal{D}_{\text{opt}}$ is responsible for transitions between excitation number manifolds, through which sunlight photons pump our superabsorber. Conversely, $\mathcal{D}_{\text{vib}}$, leads to vibrational relaxation within excitation manifolds, and, dependent on eigenstructuring, manifesting as either guide-sliding or dark-state trapping.
To quantify the efficiency of the guide-slide superabsorber, we employ a trap model for extracting and converting ring excitons to useful energy [53, 54, 55]. This is achieved by incoherently coupling an additional two-level system to the ring as an abstract load [53, 51] (see SM). Physically, the trap could be another dipole displaced below the centre of the ring with coupling to the ring dominated by second-order photon field effects , i.e. as a FRET process. From this trap dipole, excitations can then be converted to useful energy.
Combining high-temperature solar photons with room-temperature phonons and the trap, we have thus constructed an effective quantum heat engine (QHE), allowing us to determine the power output of our ring antennas, and uncover how this power scales with the number of dipoles. We utilise a standard approach for calculating the power outputs of such a QHE [53] by optimising the trap’s power with respect to its decay rate that represents the load of the device (see SM for full details).
We directly compare the power outputs of the solar absorbers using the tilted guide-slide developed previously and the novel surface plasmon-enabled (parallel) guide-slide developed here. Fig. 4 shows the results of numerical simulations for varying numbers of dipoles in the rings, for a permittivity of $\varepsilon=-2.37$, similar to that of chromium at $1.8$eV. Fig. 4c depicts the power output per dipole, and we note an increase with $N$ for the spherical guide-slide setup, showing super-extensive scaling. We see for small values of $N$ the power output achieved by the tilted guide-slide is greater than that of the parallel guide-slide. This is due to the relative size of the sphere to the ring radius being small in this limit, restricting the ability to engineer dipolar interactions. By contrast, increasing ring radius improves the approximation of dipoles experiencing a planar dielectric interface, meaning the design is scalable to larger $N$. In Fig. 4b, a log-log plot shows the exponent for each configuration’s power output of the form $P=\alpha N^{m}$. Our surface plasmon-enabled guide-slide has greater super-extensive power scaling than the tilted guide-slide, with a growth exponent $m=1.55$ vs. $m=1.08$. These growth exponents were calculated using linear regression with sum squared residuals of $0.0049$ and $0.0021$ respectively. This increase in super-linear scaling is due to the parallel dipoles in the surface plasmon-enabled guide-slide offering maximal collective enhancement.
Our results exemplify the utility of selectively designing the electromagnetic field through metals or dielectrics to achieve bespoke couplings different from far-field behaviour. While incorporating a spherical metallic nano-particle into the centre of a nanoscopic molecular ring of dipoles presents an implementation challenge, recent experimental studies have validated the approach of utilising DNA origami to interface molecular rings with nano-particles  [56, 57, 58], supporting the in-principle feasibility of our outlined design. Its simplicity, enhanced super-linear scaling, and operability without a photonically engineered environment and active re-initialisation establish our present design as a leading candidate for a quantum-enhanced molecular antenna. Structuring the local electromagnetic field modes to induce the engineered dipole-dipole interactions also impacts the optical decay rates and yields a secondary advantage (see SM). Specifically, we observe an additional enhancement of the super-linear behaviour from a $N$-dependent Purcell effect. As this effect is not directly related to realising superabsorption, as the counterpart to Dicke superradiance, to avoid ambiguity we have opted not to include this effect in this discussion. While our design—unlike previous superabsorbing ring proposals—performs well without a photonic band-gap and active reinitisalition, it benefits and maintains its advantage when available leading to greater super-extensive scaling, approaching the theoretical limit (see SM). Taking into account the effects of dephasing, detuning and absorption into the dielectric we also determined that the surface plasmon-enabled guide-slide is remarkably robust all these types of deleterious effects even at moderate levels (see SM).
[…]
Figure 4: Scaling of the solar cell power output for the configuration of the parallel spherical ring guide-slide, and the previously developed tilted guide-slide. The (a) power output (b) log-log plot of the power output (c) power output per dipole of the ring for increasing number of dipoles in the ring. The theoretical maximum Dicke scaling of $N^{2}$ (green dashed) and linear scaling (black dashed) are also plotted for comparison in (c). Full simulation parameters are given in the SM.
In summary, we have demonstrated that surface plasmons allow the engineering of molecular dipole-dipole interactions, thereby enabling advantageous couplings for light-harvesting systems. Specifically, we have shown the possibility of combining desirable far-field H-aggregate-like properties with effective J-aggregate-like near-field couplings. Building upon this, we have developed a novel superabsorber by precisely engineering the Hamiltonian of a molecular dipole ring through the strategic placement of a metallic nanoparticle at its center. Within the framework of a quantum heat engine, we have shown that our design surpasses previous proposals in performance, and, due to its greater simplicity, represents an ideal candidate for the development of light-harvesting superabsorbers. Importantly, we have illustrated that the innovative design of the local dielectric or metallic environment enables the engineering of system Hamiltonians with properties conducive to the design and implementation of quantum-enhanced devices, representing a critical advancement towards realising efficient energy transport systems. The paradigm introduced offers profound and wide-reaching impact across a broad range of quantum technologies which rely on both engineered interactions and coupling to an external environment.
Acknowledgments
The authors thank Will Brown for helpful discussions. A. Burgess and E.M. Gauger thank the Leverhulme Trust for support through grant number RPG-2022-335. M.C. Waller thanks the EPSRC Doctoral Training Programme EPSRC/DTP 2020/21/EP/T517896/1, and R. Bennett acknowledges financial support from EPSRC grant EP/W016486/1.